# Improved Performances and Motivation in Intelligent Tutoring Systems: Combining Machine Learning and Learner Choice

### v1

@@ Abstract
Although the benefits of personalized learning are now well documented, its generalization in schools is challenged by high class sizes. Educational technologies such as intelligent tutoring systems (ITS) could contribute to address this challenge and help teachers and students. Recently, an approach leveraging models of curiosity-driven learning was leveraged to build an ITS personalizing sequences of exercices. Based on the Learning Progress hypothesis (LPH), this approach consisted in proposing students exercices that maximize learning progress, and incrementally identified using multi-armed bandit machine learning techniques. The resulting algorithm (ZPDES) was previously shown in field studies to be more efficient in terms of learning performances for a diversity of student profiles when compared to a curriculum designed by a human expert. However, there were two limits. First, the motivational impact was not assessed. Second, the ZPDES algorithm did not leave students the ability to express choices. This limitation in agency was at odds with the LPH theory initially concerned with modeling curiosity-driven learning. Here, we introduce a system (ZCO) that combines adaptive exercise recommentations using LP, with the possibility for students to make choices. These choice possibilities concern dimensions that are orthogonal to exercice difficulty, and are an instance of gamification also used in many existing educational technologies.
We present an extensive field study (265 7-8 years old children from 11 schools, RCT design) comparing an LP-based automatic curriculum generation system with a hand-designed curriculum, both with and without self-choice. We first show that LP-based personalization improves learning performance (reproducing and solidifying previous results) while producing a positive and motivating learning experience. We then show that the addition of self-choice as a playful feature triggers intrinsic motivation in the learner and reinforces the learning effectiveness of the LP-based personalizing. In doing so, it strengthens the links between intrinsic motivation and performance progress during the serious game. Conversely, deleterious effects of the playful feature are observed for hand-designed linear paths. Thus, the intrinsic motivation elicited by a playful feature is beneficial only if the curriculum personalization is effective for the learner. Such a result deserves great attention due to the increased use of playful features in non adaptive educational technologies available in the market.
Keywords Adaptive Learning  $\cdot$ Learning Progress  $\cdot$ Intelligent Tutoring System  $\cdot$ Machine Learning  $\cdot$ Autodetermination theory  $\cdot$ Zone of Proximal development  $\cdot$ Flow Theory
@@ 1 Introduction
The evidence-based assets of personalized learning over one-size-fits-all educational approaches are today well documented [2, 3]. As classroom sizes are still high, it is difficult for teachers to set up individualized teaching paths, which is why high expectations are placed on Educational Technologies (ET) to automate them and support teachers in their missions.
The exploitation of artificial intelligence and digital systems has then become a crucial question to improve ET [4] and enable forms of adaptivity and personalization, leading to the development of Intelligent Tutoring Systems (ITS). This aims to make education more effective and accessible for the large diversity of students and as a way to provide useful objective metrics on learning [5, 6, 7].
Among the ITS field, there has been several approaches to optimize, personalize and adapt ITS to learners in order to enhance access to quality learning experience for all learners. Adaptation in learning technologies can be described as a structure with three main components [8].Firstly, adaptation to the instruction source, i.e. to what it will be adapted, such as the learner learning style [9, 10], knowledge [11] or preferences [12]. Secondly, the target of the adaptive instruction, i.e. what will be adapted, such as the content [9] or the presentation [13]. Thirdly, the adaptive component generating a pathway between the two first components, i.e. how to adapt a Target to a Source, such as rule-based systems [9] or Bayesian-networks [14].
This last component is the engine generating a curriculum of training activities for learners in ITS. Thus, to be able to adapt the content to the learner, all ITS are most often implicitly or explicitly based on the concepts of the zone of proximal development (ZPD) [15] and the state of Flow [16]. These are well-known concepts in developmental psychology, and have inspired many models of learning for education (e.g., Cognitive load theory[17]). Following them, many ITS aim to offer the learner pedagogical activities that are neither too difficult nor too easy with regard to their abilities, so that they can be engaged and progress in their acquisitions without being anxious or bored during the process. ITS can also propose activities the learner can not solve alone but will be able to solve with hints or with the teacher’s help. From it, the ITS can be divided in two main design categories for managing learning curricula [1] : The former, namely the "linear design" involves that all learners follow a single path although at different speeds or with a different number of attempts. The latter, called "branched-paths design", enables each learner to follow a specific path according to their their own need.
@@
In this context, the ZPDES algorithm (Zone of Proximal Development and Empirical Success) has been proposed to be used as a new LP-based activity manager in ITS[31].
@@
Leveraging multi-armed bandit algorithms[32, 33], it integrates an expert knowledge rule-based system and exploits the LP to select the activities which present the highest learning value for the student progress. Basically, ZPDES algorithm uses the student success rate to dynamically computes the best activity set (ZPD) and ponders the activities from LP to select them stochastically. The general idea is illustrated in figure 1 and shows the evolution of the possible activities available to the student as compared to what happens with a predefined linear sequence (Predef). ZPDES has been evaluated as an efficient algorithm for adaptive generation of various learning paths for real students[34], robust to heterogeneous populations as shown by complementary systematic experiments with simulated learners[35].
@@
Figure 1: The space of available activities always contain only one activity in the predefined linear sequence (Predef) while the space expends over time with ZPDES to allow a diversity of exploration and find the best activities for the learner.
@@
In summary, the automated personalization performed by ZPDES aims to lead learners into activities providing maximal learning progress, resulting in individualized learning paths. Yet, this approach had so far two major limitations. First, while its associated objective was to enhance intrinsic motivation, the motivational impact of ZPDES was not studies so far on human learners. Second, ZPDES does not lead the learner to actively make decisions related to the learning path although the theoretical LP model[19] encompasses self-decision making. Indeed, allowing self-decisions can boost the sense of agency and have a positive motivational impact, and be an efficient vector of performance [36, 37, 38]. One original argument for this design choice for ZPDES relied on findings in educational psychology, in particular those related to the self-regulation model of decision making[39] revealing that decision making can be biased and error prone, particularly in children[40] which were a priority target of this approach. A child can lack a given resource required to make an adaptive decision (e.g., lacking adequate knowledge) or some factors could constrain the person’s ability to carry out decision-making processes (e.g., under-/over-estimation of learner regarding his learning progress)[39, 40].
@@
A simple way to overcome potential decision making failures, while promoting the learner’s intrinsic motivation to perform the activity, is to combine the use of ZPDES (to control the curriculum difficulty and variety) with offering choice possibilities limited to dimensions that are orthogonal to the learning complexity (e.g. here, choice of visual objects on which to do a math exercise as shown in Fig. 1(b)): this is the principle of ZCO (Zpdes with ChOice) system introduced in this paper. Indeed, the subjective value of a task influences academic performance[41, 42] and allowing young students (3rd through 9th-grade students[43]) to express their interest or preferences stimulates their intrinsic motivation and thus could be a booster for ITS effectiveness. An open question we address here is to understand the relative contributions and interactions of LP-based curriculum personalization and choice over both learning efficiency and motivation.
@@
This RCT, approved by Inria COERLE ethical committee, involved 265 children, from 24 classes of 11 primary schools of the Bordeaux school district. The software on which children studied during this RCT, named Kidlearn ITS[34], was designed and developed specifically for this kind of experiment. This ITS aims to teach basic mathematics for children aged 7 years old through manipulation of money bills and coins (number decomposition, addition, subtraction of integers and decimals), and has been aligned to official pedagogical objectives on this topic in the national French education system. Figure 1(a) shows Kidlearn interface on a tablet the students use. They either play the role of the client or the merchant and need to compose the correct amount of money to either pay or give the change by dragging and dropping the bills and coins on the left side. You can refer to section 5.5 for more details about the pedagogical scenario and interface description. We compared 4 versions of the KidLearn ITS[34]; the Predefined sequence without (Predef) or with (PCO) learner decision, and ZPDES condition without (ZPDES) or with (ZCO) learner decision.
@@ 2 Results
The data presented here involves 265 children (Predef: 62, PCO: 59, ZPDES: 76, ZCO: 68) from schools of the Bordeaux school district. The experiment consisted of four sessions per class over two successive week (two sessions per week with at least one day break between each sessions). During the sessions, students interact with a tablet and either answer questionnaires (pre/post test, motivation questionnaires, etc) or work on Kidlearn ITS for 30 minutes (without planned interruption). The detailed experimental schedule and overall setup are presented in section 5.6.
To study the impact of LP-based personalizing and self-decision-making on student’s learning performance and motivation, we use ZPDES as the LP-based algorithm (Sec. 5.2) and a predefined sequence following a "linear design" (called "Predef" described in Sec. 5.3) as a baseline. This predefined sequence is implemented as a series of activities in which the student must have 75% success over 4 activities of the same type to pass to the next activity type. The impact of self-decision making, an assumed in the LP-model, is also studied. It takes the form of a contextual choice given to the student consisting in choosing the visual objects presented during the exercise, thus the choice does not impact the difficulty of the exercise (Sec. 5.4). This leads to the comparison of four experimental conditions, two conditions without self-decision-making: ZPDES and Predef; two conditions with it: ZCO (ZPDES with Choice of Object) and PCO (Predef with Choice of Object).
We first check the differences in curricula, i.e the student progression in the Kidlearn app scenario, between each conditions. To grasp these differences, we compare the learning activities they reach and achieve during training as well as activity space evolution through time (Sec. 2.1). Differences in curricula raise the question of the learning effectiveness of each conditions (Sec. 2.2 and their impact on learning experience and motivation (Sec. 2.3). The learning effectiveness is evaluated through comparison between pre- and post-test results, while an emotional scale is used to assess the emotional valence of learning experience and the motivation is evaluated through Vallerand’s questionnaire[50, 51] (see Sec. 5.6.3 for more information on measures). Then, the relations between LP-based personalization and subsequent learning performance and motivation are analysed (Sec. 2.4). Finally, we check if individual characteristics modulate the impact of LP-based personalization of (Sec. 2.5).
@@ 2.1 How do LP-based personalized curricula differ from hand-designed ones?
It is not always trivial to grasp meaningful statistical differences between curricula over a population of student. Thereby, in order to be able to compare globally and quantitatively the curricula generated in each conditions, an "Activity score" has been designed (defined in Sec. 5.6.3). This score represents the level of difficulty either reached or achieved (i.e. reached and succeeded) by the students for each type of activity. In other words, we use this score as a proxy to understand how the student activity space generally evolves across time for each conditions.
@@
Also, giving a contextual choice does not seem to affect the activities done by the students due to the fact that ZPDES and ZCO scores are really similar, as it is for Predef and PCO.
To have a more qualitative overview of the student activities, figure 4 shows the curriculum of each student through the activities made at $4$ different times steps. For a given time step $t$ and condition, a matrix slot represents the state of an activity (ordinate) for a particular student (abscissa). A slot is grey if a student has never explored the corresponding activity and it is purple if the student is doing this activity at time $t$. When a student has explored an activity, the slot is tinted green depending on the student’s success rate (light green: low, dark green: high).
@@
The learning effectiveness of each condition is evaluated through comparison between pre- and post-test results. The pre- and post-test (precisions in Sec. 2.2) are composed of 20 items scoring from 0 to 1 (max score is 20). The pre-test happens at the beginning of the first session, while the post-test happens at the end of the last session. The pre- and post-test are presented on the tablet on a dedicated interface (different from the ITS one). Each item of the test evaluates the student over knowledge and skills related to money manipulation, number composition, addition or subtraction (similar to the skills and knowledge trained in the ITS). Both tests include the same items organised in the same order but the items’ wording have randomly selected values for each item and each student (with verification that no items in the post-test have the same values in their wording as the ones in the pre-test for one student).
The statistical procedure used in this section consists of three-way mixed ANOVA (algo x choice x pre/post) on the Math-tests score (pre- and post-measurement of student performance), with the pre/post factor as within-subject factor.
@@
The statistical procedure used here consists of a two-way ANOVA (algo x choice) on the Emotional Scale Score (summative score of emotional valence of learning experience collected during each Kidlearn session, see Kidlearn-related learning experience in section 2.3).
@@
In order to establish a relationship between the learning effectiveness and the learning experience according to each experimental condition, correlations for each experimental condition (Predefined, PCO, ZPDES, and ZCO) were made between the following 3 measures (see Tab. 1) : 1) Learning score (difference between pre-and post-test) ; 2) Kidlearn progression (Final activity score defined in section 5.6.3) ; 3) Motivation score (see Sec. 5.6.3).
@@ 2.5 Is the impact of LP-based personalization modulated by individual characteristics of learners?
All the previous analyses have in addition been conducted with ANCOVA analyses where the covariable was related to individual characteristics. Particularly we investigated the mediating effect of school satisfaction, digital technology experience, gender and age (see Profil metrics in section 5.6.3).
No significant results have been observed revealing that the present results are robust to mediating effects related to the studied individual factors (see table 4 in appendix).

### v2

@@ Abstract
Large class sizes challenge personalized learning in schools, prompting the use of educational technologies such as intelligent tutoring systems. To address this, we present an AI-driven personalization system, called ZPDES, based on the Learning Progress Hypothesis - modeling curiosity-driven learning - and multi-armed bandit techniques. It sequences exercises that maximize learning progress for each student. While previous studies demonstrated its efficacy in enhancing learning compared to hand-made curricula, its impact on student motivation remained unexplored. Furthermore, ZPDES previously lacked features allowing student choice, a limitation in agency that conflicts with its foundation on models of curiosity-driven learning.
This study investigates how integrating choice, as a gamification element unrelated to exercise difficulty, affects both learning outcomes and motivation. We conducted an extensive field study (265 7-8 years old children, RCT design), comparing ZPDES with and without choice against a hand-designed curriculum. Results show that ZPDES improves both learning performance and the learning experience. Moreover adding choice to ZPDES enhances intrinsic motivation and further strengthens its learning benefits. In contrast, incorporating choice into a fixed, linear curriculum negatively impacts learning outcomes.
These findings highlight that the intrinsic motivation elicited by choice (gamification) is beneficial only when paired with an adaptive personalized learning system. This insight is critical as gamified features become increasingly prevalent in educational technologies.
Keywords Educational technologies  $\cdot$ Artificial Intelligenc  $\cdot$ Adaptive Learning  $\cdot$ Intelligent Tutoring System  $\cdot$ Intrinsic Motivation  $\cdot$ Gamification
@@ 1 Introduction
The evidence-based assets of personalized learning over one-size-fits-all educational approaches are today well documented [2, 3]. As classroom sizes are still high, it is difficult for teachers to set up individualized teaching paths, which is why high expectations are placed on Educational Technologies (ET) to automate them and support teachers in their missions. The exploitation of artificial intelligence and digital systems has then become a crucial question to improve ET [4] and enable forms of adaptivity and personalization, leading to the development of Intelligent Tutoring Systems (ITS). This aims to make education more effective and accessible for the large diversity of students and as a way to provide useful objective metrics on learning [5, 6, 7].
Among the ITS field, there has been several approaches to optimize, personalize and adapt ITS to learners in order to enhance access to quality learning experience for all learners. Adaptation in learning technologies can be described as a structure with three main components [8]. Firstly, adaptation to the instruction source, i.e. to what it will be adapted, such as the learner learning style [9, 10], knowledge [11] or preferences [12]. Secondly, the target of the adaptive instruction, i.e. what will be adapted, such as the content [9] or the presentation [13]. Thirdly, the adaptive component generating a pathway between the two first components, i.e. how to adapt a Target to a Source, such as rule-based systems [9] or Bayesian-networks [14]. This last component is the engine generating a curriculum of training activities for learners in ITS. Thus, to be able to adapt the content to the learner, all ITS are most often implicitly or explicitly based on the concepts of the zone of proximal development (ZPD) [15] and the state of Flow [16].
These are well-known concepts in developmental psychology, and have inspired many models of learning for education (e.g., Cognitive load theory[17]). Following them, many ITS aim to offer the learner pedagogical activities that are neither too difficult nor too easy with regard to their abilities, so that they can be engaged and progress in their acquisitions without being anxious or bored during the process. ITS can also propose activities the learner can not solve alone but will be able to solve with hints or with the teacher’s help. From it, the ITS can be divided in two main design categories for managing learning curricula [1]. The former, namely the "linear design" involves that all learners follow a single path although at different speeds or with a different number of attempts. The latter, called "branched-paths design", enables each learner to follow a specific path according to their their own need.
@@
In this context, the ZPDES algorithm (Zone of Proximal Development and Empirical Success) has been proposed to be used as a new LP-based activity manager in ITS[31]. Leveraging multi-armed bandit algorithms[32, 33], it integrates an expert knowledge rule-based system and exploits the LP to select the activities which present the highest learning value for the student progress. Basically, ZPDES algorithm uses the student success rate to dynamically computes the best activity set (ZPD) and ponders the activities from LP to select them stochastically. The general idea is illustrated in figure 1 and shows the evolution of the possible activities available to the student as compared to what happens with a predefined linear sequence (Predef). ZPDES has been evaluated as an efficient algorithm for adaptive generation of various learning paths for real students[34], robust to heterogeneous populations as shown by complementary systematic experiments with simulated learners[35].
@@
Figure 1: The space of available activities always contain only one activity in the predefined linear sequence (Predef) while the space expends over time with ZPDES to allow a diversity of exploration and find the best activities for the learner. The type of activity is represented by a letter (A or B), and the difficulty by a number (1,2 or 3).
@@
In summary, the automated personalization performed by ZPDES aims to lead learners into activities providing maximal learning progress, resulting in individualized learning paths. Yet, this approach had so far two major limitations. First, while its associated objective was to enhance intrinsic motivation, the motivational impact of ZPDES was not studies so far on human learners. Second, ZPDES does not lead the learner to actively make decisions related to the learning path although the theoretical LP model[19] encompasses self-decision making.
Indeed, allowing self-decisions can boost the sense of agency and have a positive motivational impact, and be an efficient vector of performance [36, 37, 38, 39]. One original argument for this design choice for ZPDES relied on findings in educational psychology, in particular those related to the self-regulation model of decision making[40] revealing that decision making can be biased and error prone, particularly in children[41] which were a priority target of this approach. A child can lack a given resource required to make an adaptive decision (e.g., lacking adequate knowledge) or some factors could constrain the person’s ability to carry out decision-making processes (e.g., under-/over-estimation of learner regarding his learning progress)[40, 41].
@@
A simple way to overcome potential decision making failures, while promoting the learner’s intrinsic motivation to perform the activity, is to combine the use of ZPDES (to control the curriculum difficulty and variety) with offering choice possibilities limited to dimensions that are orthogonal to the learning complexity (e.g. here, choice of visual objects on which to do a math exercise as shown in Fig. 2(b)). This is the principle of ZCO [^1] ZCO stands for Zone of proximal development with ChOice system introduced in this paper. Indeed, the subjective value of a task influences academic performance[42, 43] and allowing young students (3rd through 9th-grade students[44]) to express their interest or preferences stimulates their intrinsic motivation and thus could be a booster for ITS effectiveness. An open question we address here is to understand the relative contributions and interactions of LP-based curriculum personalization and choice over both learning efficiency and motivation.
@@
This RCT, approved by Inria COERLE ethical committee, involved 265 children, from 24 classes of 11 primary schools of the Bordeaux school district. The software on which children studied during this RCT, named Kidlearn ITS[34], was designed and developed specifically for this kind of experiment. This ITS aims to teach basic mathematics for children aged 7 years old through manipulation of money bills and coins (number decomposition, addition, subtraction of integers and decimals), and has been aligned to official pedagogical objectives on this topic in the national French education system. Figure 2(a) shows Kidlearn interface on a tablet the students use. They either play the role of the client or the merchant and need to compose the correct amount of money to either pay or give the change by dragging and dropping the bills and coins on the left side. You can refer to 5.2 for more details about the pedagogical scenario and interface description. We compared 4 versions of the KidLearn ITS[34]; the Predefined sequence without (Predef) or with (PCO) learner decision, and ZPDES condition without (ZPDES) or with (ZCO) learner decision.
@@ 2 Results
The data presented here involves 265 children (Predef: 62, PCO: 59, ZPDES: 76, ZCO: 68) from schools of the Bordeaux school district. The experiment consisted of four sessions per class over two successive week (two sessions per week with at least one day break between each sessions). During the sessions, students interact with a tablet and either answer questionnaires (pre/post test, motivation questionnaires, etc) or work on Kidlearn ITS for 30 minutes (without planned interruption). The detailed experimental schedule and overall setup are presented in 5.3.
To study the impact of LP-based personalizing and self-decision-making on student’s learning performance and motivation, we use ZPDES as the LP-based algorithm (see 5.1.2) and a predefined sequence following a "linear design" (called "Predef" described in 5.1.3) as a baseline. This predefined sequence is implemented as a series of activities in which the student must have 75% success over 4 activities of the same type to pass to the next activity type. The impact of self-decision making, an assumed in the LP-model, is also studied. It takes the form of a contextual choice given to the student consisting in choosing the visual objects presented during the exercise, thus the choice does not impact the difficulty of the exercise (details in 5.1.4). This leads to the comparison of four experimental conditions, two conditions without self-decision-making: ZPDES and Predef; two conditions with it: ZCO (ZPDES with Choice of Object) and PCO (Predef with Choice of Object).
We first check the differences in curricula, i.e the student progression in the Kidlearn app scenario, between each conditions. To grasp these differences, we compare the learning activities they reach and achieve during training as well as activity space evolution through time. Differences in curricula raise the question of the learning effectiveness of each conditions and their impact on learning experience and motivation. The learning effectiveness is evaluated through comparison between pre- and post-test results, while an emotional scale is used to assess the emotional valence of learning experience and the motivation is evaluated through Vallerand’s questionnaire[51, 52] (see 5.3.3 for more information on measures). Then, the relations between LP-based personalization and subsequent learning performance and motivation are analysed.Finally, we check if individual characteristics modulate the impact of LP-based personalization.
@@ 2.1 How do LP-based personalized curricula differ from hand-designed ones?
It is not always trivial to grasp meaningful statistical differences between curricula over a population of student. Thereby, in order to be able to compare globally and quantitatively the curricula generated in each conditions, an “Activity score” has been designed (details in 5.3.3). This score represents the level of difficulty either reached or achieved (i.e. reached and succeeded) by the students for each type of activity. In other words, we use this score as a proxy to understand how the student activity space generally evolves across time for each conditions.
@@
Also, giving a contextual choice does not seem to affect the activities done by the students due to the fact that ZPDES and ZCO scores are really similar, as it is for Predef and PCO. To have a more qualitative overview of the student activities, figure 4 shows the curriculum of each student through the activities made at $4$ different times steps. For a given time step $t$ and condition, a matrix slot represents the state of an activity (ordinate) for a particular student (abscissa). A slot is grey if a student has never explored the corresponding activity and it is purple if the student is doing this activity at time $t$. When a student has explored an activity, the slot is tinted green depending on the student’s success rate (light green: low, dark green: high).
@@
The learning effectiveness of each condition is evaluated through comparison between pre- and post-test results. The pre- and post-test are composed of 20 items scoring from 0 to 1 (max score is 20). The pre-test happens at the beginning of the first session, while the post-test happens at the end of the last session. The pre- and post-test are presented on the tablet on a dedicated interface (different from the ITS one). Each item of the test evaluates the student over knowledge and skills related to money manipulation, number composition, addition or subtraction (similar to the skills and knowledge trained in the ITS). Both tests include the same items organised in the same order but the items’ wording have randomly selected values for each item and each student (with verification that no items in the post-test have the same values in their wording as the ones in the pre-test for one student).
The statistical procedure used in this section consists of three-way mixed ANOVA (algo x choice x pre/post) on the Math-tests score (pre- and post-measurement of student performance), with the pre/post factor as within-subject factor. The algorithm factor includes the two conditions (ZPDES or Predef). And the choice factor include also two conditions (with or without choice). The p-value threshold is $\alpha=0.05$. Pairwise comparisons are carried out with the Least Significant Difference (LSD) and Bonferroni procedure for corrected comparisons.
@@
The statistical procedure used here consists of a two-way ANOVA (algo x choice) on the Emotional Scale Score (summative score of emotional valence of learning experience collected during each Kidlearn session, see Kidlearn-related learning experience section).
@@
In order to establish a relationship between the learning effectiveness and the learning experience according to each experimental condition, correlations for each experimental condition (Predefined, PCO, ZPDES, and ZCO) were made between the following 3 measures (see Tab. 1) : 1) Learning score (difference between pre-and post-test) ; 2) Kidlearn progression (final activity score defined in 5.3.3) ; 3) Motivation score (defined in 3).
@@ 2.5 Is the impact of LP-based personalization modulated by individual characteristics of learners ?
All the previous analyses have in addition been conducted with ANCOVA analyses where the covariable was related to individual characteristics. Particularly we investigated the mediating effect of school satisfaction, digital technology experience, gender and age (see Profil metrics in 5.3.3).
No significant results have been observed revealing that the present results are robust to mediating effects related to the studied individual factors (see Tab. 5 in 5.4).